# Splitting of electronic spectrum in paramagnetic phase of itinerant ferromagnets and altermagnets

### v1

@@ Abstract
We study self-energy effects, induced by strong magnetic fluctuations in paramagnetic phase of strongly-correlated itinerant magnets within the density functional theory combined with the dynamical mean field theory (DFT+DMFT approach) and its non-local extension. As concrete examples, we consider $\alpha$-iron, half metal CrO2, van der Waals material CrTe2, and altermagnet CrSb. We show that both local and non-local magnetic correlations yield splitting of the electronic spectrum in the paramagnetic phase, such that it closely resembles the DFT band structure in the ordered phase and suppresses spectral weight at the Fermi level. The relative importance of non-local vs. local correlations depends on the proximity to half filling of $d$ states: closer to half filling, the role of local correlations increases. Although the obtained split bands do not possess a certain spin projection, their splitting suppresses spectral weight at the Fermi level. The obtained electronic states are also expected to be easily spin polarized by a weak external magnetic field.
@@
The situation is different in the paramagnetic phase even sufficiently close to the magnetic transition. While in two-dimensional systems the so-called quasi-splitting of Fermi surfaces may occur QS1 ; QS2 ; QS3 , in three-dimensional systems the magnetic correlations are generally weaker, and not sufficient to establish a pronounced change of electronic spectrum. At the same time, two effects appear to be important in the paramagnetic phase. The first effect is the presence of local correlations, which originate mainly from Hund interaction in the so-called Hund metals Hund1 ; Hund2 ; Hund3 ; Hund4 . They lead to the formation of local magnetic moments and non-Fermi-liquid behavior and/or non-quasiparticle form of some of the electronic bands. Notably, many “classical” magnets, e.g. iron OurFe ; Sangiovanni , as well as many other transition metals and their compounds, belong to the class of Hund metals. Although the importance of this effect for quasi-splitting of the Fermi surfaces was discussed previously for the two-orbital model QS2 , the study of its effect in multi-orbital is of certain interest.
@@
General formalism. We consider the DFT+DMFT approach with the tight-binding multiorbital Hubbard model
@@
Numerical results. For numerical analysis we consider $\alpha$-iron, CrO2, CrTe2 and altermagnet CrSb. For DFT calculations we use the Quantum Espresso package QE with ultrasoft pseudopotentials from the SSSP PBEsol Precision library PPP , with subsequent Wannier projection using maximally localized Wannier functions, obtained within the Wannier90 package Wannier90 . The DFT+DMFT approach to study exchange interactions in the three former metallic ferromagnets was considered in Refs. OurJq ; MyCrO2 ; CrTe2 . We use the Slater parametrization of Coulomb interaction with around-mean-field form of the double counting correction AMF . For DMFT calculations we use segment continuous time Monte-Carlo (CT-QMC) CT-QMC iQIST solver iQIST , which allows us to considerably reduce the stochastic noise and obtain the details of the frequency dependence of the vertex functions over a sufficiently broad frequency range. We also use the corrections for finite frequency box in Bethe-Salpeter equations My_BS ; further details can be found in Refs. OurJq ; MyCrO2 ; CrTe2 . The obtained Green’s functions, containing local or non-local self-energies, are analytically continued to the real axis using ana_cont package AnaCont .
@@
In Fig. 1 we compare the DFT band structures of iron in non-magnetic and ferromagnetic phases with the spectral density in DFT+DMFT approach and the result, accounting for the non-local contributions to the self-energy in Eq. (3). The DMFT approach is applied at $\beta=1/(k_{B}T)=4.6$ eV-1 ($T=2900$ K) at $U=F^{0}=4$ eV and Hund exchange $J=0.9$ eV. Relatively high temperature is chosen in view of high Curie temperature of iron in the density-density approach, which corresponds to $\beta_{c}=4.87$ eV-1($T_{C}=2380$ K). The respective uniform susceptibility at the considered temperature $\chi_{{\mathbf{q}}=0}=105$ eV-1 is much larger than the local transverse susceptibility $\chi_{\rm loc}=12$ eV-1. One can see (Fig. 1a) that the spectral density in DMFT approach is sufficiently close to the DFT band structure in the non-magnetic phase, with a slight shift of the flat band in the direction $H$-$P$-NN slightly away from the Fermi level. The non-local effects (see Fig. 1b) yield a stronger shift of this band and a resulting qualitative change of the band structure, which becomes remarkably similar to the DFT band structure in the ferromagnetic phase due to the non-local spin fluctuations. In particular, we observe splitting of the band structure along the $H$-$P$-NN direction in momentum space (where the extended van Hove singularity was present). This splitting is reminiscent of the earlier discussed quasi-splitting in the single-band two-dimensional systems QS1 ; QS2 ; QS3 , but it differs from that case by both dimensionality and multi-band electronic structure. We note that the obtained features of electronic structure persist in a wider temperature range, with a continuous shift towards the non-magnetic band structure, see Supplemental Material.
@@
In Fig. 2 we plot the spectral functions of DFT+DMFT and non-local approach to CrO2 at $\beta=6.5$ eV-1 in comparison with the DFT band structures for nonmagnetic and ferromagnetic phases. We use the interaction strength considered previously in Refs. MyCrO2 ; Solovyev_rev , $F^{0}=1.99$ eV, $F^{2}=7.67$ eV, and $F^{4}=5.48$ eV. In the Wannier projection, we include both Cr $d$ and oxygen $p$ states (which results in an 11-orbital model per Cr site, see Ref. MyCrO2 ), but we have verified that qualitatively similar results are obtained within 5-orbital model, including Cr $d$ states only. The considered temperature is also chosen to be sufficiently close to the DFT+DMFT Curie temperature $\beta_{c}=6.8$ eV-1 ($T_{C}=1700$ K). The respective uniform susceptibility at the considered temperature $\chi_{{\mathbf{q}}=0}=110$ eV-1 is also much larger than the local susceptibility $\chi_{\rm loc}=12$ eV-1, and they are qualitatively close to those considered for iron. In case of CrO2, DMFT yields somewhat stronger change of electronic band structure than in iron; in particular it shifts the flat band in the direction $Z$-$R$-AA away from the Fermi level, which makes the band structure similar to that obtained within DFT for the ordered phase. Considering the effect of the non-local self-energy yields stronger splitting of bands, yielding again qualitative agreement with the ferromagnetic DFT band structure. Stronger effect of the local self-energy in CrO2 in comparison to iron can be attributed to somewhat closer proximity to half filling of $d$ stated ($n_{d}^{{\rm CrO}_{2}}=3.76$ vs. $n_{d}^{\rm Fe}=6.4$).
@@
In Fig. 3 we present the results for bulk van der Waals magnet CrTe2 at $T=1600$ K ($\beta=7.3$ eV-1) which is above the magnetic transition temperature $T_{c}=950$ K ($\beta_{c}=12.2$ eV-1; we choose $U=2.8$ eV, $J=0.9$ eV). We obtain the maximum of the susceptibility $\chi_{{\bf q}_{A}}=95$ eV-1, which corresponds to the tendency of antiferromagnetic ordering between the layers CrTe2 , and $\chi_{\rm loc}=41$ eV-1. In that case, already in DFT+DMFT approach, we find strong band splitting, which originates from the local magnetic correlations. This can be attributed to the relatively large local susceptibility, related to the proximity to half filling ($n_{d}=4.8$). The effect of the non-local self-energy (not shown) is sufficiently small in this case, and does not change qualitatively (and weakly changes quantitatively) the resulting band structure (we have verified that the non-local corrections remain weak also at $\beta=9$ eV-1 closer to the DMFT phase transition). Although CrTe2 has weak interlayer coupling and therefore represents the material for which quasi-splitting of the Fermi surface originating from magnetic correlations proposed previously for two-dimensional materials QS1 ; QS2 ; QS3 could be active, we find that the effect of strong local correlations is much more pronounced due to close proximity to half filling. In view of weaker dependence of local magnetic susceptibility, than the non-local one, on temperature, we expect that band splitting in this case can be observed in a broader temperature range. Remarkably, recently the band splitting above Curie temperature was discussed for single-layer CrTe2, deposited on bilayer graphene CrTe2exp .

### v2

@@ Abstract
We study self-energy effects, induced by strong magnetic fluctuations in paramagnetic phase of strongly-correlated itinerant magnets within the density functional theory combined with the dynamical mean field theory (DFT+DMFT approach) and its non-local extension. As concrete examples, we consider $\alpha$-iron, half-metal CrO2, van der Waals material CrTe2, and altermagnet CrSb. We show that both local and non-local magnetic correlations yield splitting of the electronic spectrum in the paramagnetic phase, such that it closely resembles the DFT band structure in the ordered phase. The relative importance of non-local vs. local correlations depends on the proximity to half filling of $d$ states: closer to half filling, the role of local correlations increases. Although the obtained split bands do not possess a certain spin projection, their splitting suppresses spectral weight at the Fermi level. The obtained electronic states are also expected to be easily spin polarized by a weak external magnetic field.
@@
The situation is different in the paramagnetic phase even sufficiently close to the magnetic transition. While in two-dimensional systems the so-called quasi-splitting of Fermi surfaces may occur QS1 ; QS2 ; QS3 ; QS4 , in three-dimensional systems the magnetic correlations are generally weaker, and not sufficient to establish a pronounced change of electronic spectrum. At the same time, two effects appear to be important in the paramagnetic phase. The first effect is the presence of local correlations, which originate mainly from Hund interaction in the so-called Hund metals Hund1 ; Hund2 ; Hund3 ; Hund4 . They lead to the formation of local magnetic moments and non-Fermi-liquid behavior and/or non-quasiparticle form of some of the electronic bands. Notably, many “classical” magnets, e.g. iron OurFe ; Sangiovanni , as well as many other transition metals and their compounds, belong to the class of Hund metals. Although the importance of this effect for quasi-splitting of the Fermi surfaces was discussed previously for the two-orbital model QS2 , the study of its effect in multi-orbital is of certain interest.
@@
General formalism. We consider the DFT+DMFT approach with the tight-binding multi-orbital Hubbard model
@@
Numerical results. For numerical analysis we consider $\alpha$-iron, CrO2, CrTe2 and altermagnet CrSb. For DFT calculations we use the Quantum Espresso package QE with ultrasoft pseudopotentials from the SSSP PBEsol Precision library PPP , with subsequent Wannier projection using maximally localized Wannier functions, obtained within the Wannier90 package Wannier90 . The DFT+DMFT approach to study exchange interactions in the three former metallic ferromagnets was considered in Refs. OurJq ; MyCrO2 ; CrTe2 . We use parametrization of Coulomb interaction by Slater integrals $F^{i}$ with around-mean-field form of the double counting correction AMF . For DMFT calculations we use segment continuous time Monte-Carlo (CT-QMC) CT-QMC iQIST solver iQIST , which allows us to considerably reduce the stochastic noise and obtain the details of the frequency dependence of the vertex functions over a sufficiently broad frequency range. We also use the corrections for finite frequency box in Bethe-Salpeter equations My_BS ; further details can be found in Refs. OurJq ; MyCrO2 ; CrTe2 . The obtained Green’s functions, containing local or non-local self-energies, are analytically continued to the real axis using ana_cont package AnaCont .
@@
In Fig. 1 we compare the DFT band structures of iron in non-magnetic and ferromagnetic phases with the spectral density in DFT+DMFT approach and the result, accounting for the non-local contributions to the self-energy in Eq. (3). The DMFT approach is applied at $\beta=1/(k_{B}T)=4.6$ eV-1 ($T=2900$ K) at $U=F^{0}=4$ eV and Hund exchange $J=(F^{2}+F^{4})/14=0.9$ eV, considered previously in Refs. OurJq ; MySU2 . Relatively high temperature is chosen in view of high Curie temperature of iron in the density-density approach, which corresponds to $\beta_{c}=4.87$ eV-1($T_{C}=2380$ K). The respective uniform susceptibility at the considered temperature $\chi_{{\mathbf{q}}=0}=105$ eV-1 is much larger than the local transverse susceptibility $\chi_{\rm loc}=12$ eV-1. One can see (Fig. 1a) that the spectral density in DMFT approach is sufficiently close to the DFT band structure in the non-magnetic phase, with a slight shift of the flat band in the direction $H$-$P$-NN slightly away from the Fermi level. The non-local effects (see Fig. 1b) yield a stronger shift of this band and a resulting qualitative change of the band structure, which becomes remarkably similar to the DFT band structure in the ferromagnetic phase due to the non-local spin fluctuations. In particular, we observe splitting of the band structure along the $H$-$P$-NN direction in momentum space (where the extended van Hove singularity was present). This splitting is reminiscent of the earlier discussed quasi-splitting in the single-band two-dimensional systems QS1 ; QS2 ; QS3 ; QS4 , but it differs from that case by both dimensionality and multi-band electronic structure. We note that the obtained features of electronic structure persist in a wider temperature range, with a continuous shift towards the non-magnetic band structure, see Supplemental Material.
@@
In Fig. 2 we plot the spectral functions of DFT+DMFT and non-local approach to CrO2 at $\beta=6.5$ eV-1 in comparison with the DFT band structures for nonmagnetic and ferromagnetic phases. We use the interaction strength considered previously in Refs. MyCrO2 ; Solovyev_rev , $F^{0}=1.99$ eV, $F^{2}=7.67$ eV, and $F^{4}=5.48$ eV. In the Wannier projection, we include both Cr $d$ and oxygen $p$ states (which results in an 11-orbital model per Cr site, see Ref. MyCrO2 ), but we have verified that qualitatively similar results are obtained within the 5-orbital model, including Cr $d$ states only. The considered temperature is also chosen to be sufficiently close to the DFT+DMFT Curie temperature $\beta_{c}=6.8$ eV-1 ($T_{C}=1700$ K). The respective uniform susceptibility at the considered temperature $\chi_{{\mathbf{q}}=0}=110$ eV-1 is also much larger than the local susceptibility $\chi_{\rm loc}=12$ eV-1, and they are qualitatively close to those considered for iron. In the case of CrO2, DMFT yields somewhat stronger change in the electronic band structure than in iron; in particular, it shifts the flat band in the direction $Z$-$R$-AA away from the Fermi level, which makes the band structure similar to that obtained within DFT for the ordered phase. Considering the effect of the non-local self-energy yields a stronger splitting of bands, yielding again qualitative agreement with the ferromagnetic DFT band structure. The stronger effect of the local self-energy in CrO2 in comparison to iron can be attributed to somewhat closer proximity to half-filling of $d$ states ($n_{d}^{{\rm CrO}_{2}}=3.76$ vs. $n_{d}^{\rm Fe}=6.4$).
@@
In Fig. 3 we present the results for bulk van der Waals magnet CrTe2 at $T=1600$ K ($\beta=7.3$ eV-1) which is above the magnetic transition temperature $T_{c}=950$ K ($\beta_{c}=12.2$ eV-1; we choose $U=F^{0}=2.8$ eV, $J=0.9$ eV, corresponding to $U_{K}=3.9$ eV, $J_{K}=0.65$ eV in Kanamori parametrization cRPA ). We obtain the maximum of the susceptibility $\chi_{{\bf q}_{A}}=95$ eV-1, which corresponds to the tendency of antiferromagnetic ordering between the layers CrTe2 , and $\chi_{\rm loc}=41$ eV-1. In that case, already in DFT+DMFT approach, we find strong band splitting, which originates from the local magnetic correlations. This can be attributed to the relatively large local susceptibility, related to the proximity to half filling ($n_{d}=4.8$). The effect of the non-local self-energy (not shown) is sufficiently small in this case, and does not change qualitatively (and weakly changes quantitatively) the resulting band structure (we have verified that the non-local corrections remain weak also at $\beta=9$ eV-1 closer to the DMFT phase transition). Although CrTe2 has weak interlayer coupling and therefore represents the material for which quasi-splitting of the Fermi surface originating from magnetic correlations proposed previously for two-dimensional materials QS1 ; QS2 ; QS3 ; QS4 could be active, we find that the effect of strong local correlations is much more pronounced due to close proximity to half filling. In view of weaker dependence of local magnetic susceptibility, than the non-local one, on temperature, we expect that band splitting in this case can be observed in a broader temperature range. Remarkably, recently the band splitting above Curie temperature was discussed for single-layer CrTe2, deposited on bilayer graphene CrTe2exp .